OpenStudy (anonymous):

Marks: 1 Simplify \frac{\sqrt[3]{7x^{5}}}{\sqrt[3]{56x^{2}}} Choose one answer. a. 2x b. \frac{x}{2} c. \frac{2}{x} d. \frac{1}{2x}

OpenStudy (anonymous):

hey, for latex, you should put \(\[\) and \(\]\) in the front and the end of the expression to make it popping up. hihihi

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

\[\frac{\sqrt[3]{7x^{5}}}{\sqrt[3]{56x^{2}}}\]

OpenStudy (anonymous):

dont know how to type it in like that

OpenStudy (anonymous):

b. \(\frac{x}{2}\) c. \(\frac{2}{x}\) d.\( \frac{1}{2x}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

ohh cause im just copying and pasting , hurrying up rynna get done with this hw . i appreciate the "how to" tutorial but im just looking for answers

OpenStudy (anonymous):

To me, you copy the problem with the options of other problem because none of them is the correct answer.

OpenStudy (anonymous):